chad drove 3 tinmes as fast as katy pedaled. In fact, chad drove 90 miles in 2 hours less than it took katy to pedal 40 miles. how fast did each of them travel? how long did they travel?
Answer by mananth(16946)

katy's speed = x
chad's speed = 3x
..
time taken by Katy for 40 miles = 40/x
time taken by Chad for 90 miles = 90/3x
40/x -90/3x=2 hours
LCM = 3x
120-90 / 3x= 2
Multiply by 3x on both sides
30*3x/3x= 3x*2
30=6x
divide by 6 on both sides
30/6=6x/6
x=5 mph Katy's speed
Time taken by Katy = 40/15= 2.66
Chad 's speed = 3x= 15 mph
Time taken by chad = 90/15=6 hours